# Stability analysis through folds: An end-loaded elastica with a lever arm

### v5

@@ 1 Introduction
The study of physical problems using the variational framework has a rich history. In this setting, the equilibria are characterized as critical points of an energy-like functional. Typically, these problems are expressed in terms of a control parameter, generating a parameter family of equilibria. Indeed, the equilibria are computed by a slow scan of a parameter from a known solution using a continuation technique [1]. However, not all equilibria are stable and physically relevant; they must correspond to the local minima of the functional. Hence, it is crucial to determine whether the critical points are local minima (or stable) and track their changes. Generally, the stability of parameter-dependent equilibria changes at singularities, such as a fold or bifurcation in the parameter space [2, 3, 4]. However, it remains unclear whether a fold in the unstable equilibrium results in a stable equilibrium or a higher-order unstable equilibrium. The direction of stability transitions can be encoded through a particular projection of solutions, using distinguished bifurcation diagrams, and the stability can be assessed with a little observation and without any additional analysis. This idea of topologically exploiting the variational structure for stability examination was first employed in Astrophysics [5, 6]. Distinguished bifurcation diagrams were later developed for finite-dimensional mechanical systems [7] that are subsequently extended to infinite-dimensional systems [8, 9]. Since then, several versions of these diagrams have appeared in investigations of elastic problems [10, 11]. In most studies, the diagrams are limited to cases with fixed-fixed ends or cases where the varying parameter appears within the integrand or as a Lagrange multiplier. In this study, we generalize the distinguished bifurcation diagrams to variational problems subject to fixed-free ends, when the bifurcation parameter appears in the boundary conditions. Consequently, two relevant cases arise: when the parameter appears at the fixed end and when it appears at the free end. For this analysis, we introduce the notion of a stability index that is equal to the dimension of the subspace over which the associated second variation operator is negative. The critical points with index zero are local minima and stable. Alternatively, the energy of the system can be used to determine the direction of stability [12, 13]. The lower branch of the fold corresponds to the lower-energy state and therefore, local minima.
@@
The developed framework is applied to test the stability of elastica under an end load through a lever arm. Cantilever structures with a lever arm are common in nature and engineering—for example, in plants supporting fruits or in load-bearing flexible systems. In recent years, nonlinear rod models [14] have been actively employed in the fields of biophysics [15], computer graphics [16, 17] and soft robot manipulators [18, 19]. Elastica, a simple planar version of this rod model, has equally received attention. A flexible elastic structure subjected to loads or constraints may exhibit multiple equilibria. Hence, it is crucial to determine their stability, as only stable structures are physically viable. Unstable equilibria may exist under specific constraints or boundary conditions and transitions to the available stable equilibrium under disturbances. Indeed, the presence of instability also hints at snap-back phenomena, a characteristic of multi-stable systems [20]. Numerous studies have been conducted on the stability of elastic rods, which involve either the Jacobi test [21, 22, 23, 24, 25, 26] or Hessian eigenvalue determination [27, 28]. Alternative techniques based on bifurcation analysis are also available [29]. In this article, we assess the stability without performing any such rigorous computations. But, instead, conclusions are drawn by qualitatively examining the plot of appropriate projection of the equilibria against the parameter.
@@
Advances in material science produced highly deformable alloys and polymers, accelerating the development of soft robotics. Generally, soft robots employ very flexible structures to generate compliant mechanisms. A flexible elastic structure under load can exhibit multiple equilibria, where one equilibrium may transition to another through snap-back instability under disturbances. These bistable or multi-stable systems exhibit large rapid displacements for small stimuli when operated around the regions of instabilities. As the system surpasses the initial energy barrier, it transitions to a new state, accompanied by a release of energy. Some of nature’s quickest mechanisms seen in Mantis shrimp [30], Venus fly trap [31], and Hummingbird beak [32] can be attributed to this snap-back instability. These mechanisms are widely used in the design of high-performance soft actuators [33], soft robots [34], deployable and morphing structures [35] and haptic devices [36]. Inspired by these expanding applications, we aim to investigate elastica with a lever arm and the dependence of its stability on system parameters. These findings could offer invaluable insights for designing innovative mechanisms. Particularly, this study signifies the distinguished bifurcation diagrams for the analysis of multi-stable systems in engineering applications.
@@ 2 The Unconstrained Variational Problem
In this section, we extend the theory of distinguished bifurcation diagrams developed by [8] and [9] to variational problems subject to fixed-free ends. We consider both cases, where the bifurcation parameter appears at the fixed end and when it appears at the free end. In section 2.1, the problem and the relevant notation are introduced. The distinguished diagram at a simple fold when the parameter appears at the fixed end is developed in section 2.2 and when the parameter appears at the free end is developed in section 2.3.
@@ 2.1 Problem Formulation
while the other end $s=l$ is held free. The scalar function $\mathcal{L}$ has continuous second derivatives with respect to all its arguments and is convex in its second argument. On the other hand, the scalar function $B$ has continuous second derivatives with respect to its first argument and has continuous first derivatives with respect to the other arguments. The system can be influenced by the parameters $\xi_{0}$ and $\xi_{l}$ at the boundaries $s=0$ and $s=l$ respectively. In the current study, we vary either $\xi_{0}$ or $\xi_{l}$, but not both at a time. Maddocks [8] generated the bifurcation diagrams for problems when the parameter appears in the integrand and with homogeneous boundary conditions, which were later extended to the case when the parameter appears in the fixed end in a problem with non-homogeneous boundary conditions [9]. However, these studies are confined to cases of fixed-fixed ends. Here, we extend the analysis to the case where one of the fixed ends is set free. The first-order condition yields the critical points $\boldsymbol{\zeta}_{o}(s,\tau,\xi_{i})$, $i=0$ or $l$, of (1) as solutions to the well-known Euler-Lagrange equations
@@
The notation $\xi_{i}$ refers to either $\xi_{0}$ or $\xi_{l}$ when not explicitly specified. These critical points $\boldsymbol{\zeta}_{o}(s,\tau,\xi_{i})$ exist as a continuous curve of solutions along a branch for some pseudo-arclength parameterization $\tau$ [1] satisfying the property
@@
The notation $\dot{\boldsymbol{\zeta}}(s,\tau)$ denotes the derivative with respect to pseudo-arclength $\tau$, whereas ${\boldsymbol{\zeta}}^{\prime}(s,\tau)$ denotes the derivative with respect to independent parameter $s$. The second variation $\delta^{2}J$ evaluated at the critical points $\boldsymbol{\zeta}_{o}(s,\tau,\xi_{i})$, must satisfy the condition [37]:
@@
and $\langle\cdot,\cdot\rangle$ represents the standard inner-product in $L^{2}$- space. The coefficient $\frac{1}{2}$ has no effect on the results and is omitted in the subsequent analysis. We define an notion of index [38] which denotes the dimension of the $\mathbf{h}$ subspace over which the second variation denoted by (9) is negative. The zero index solutions satisfy the necessary condition for local minima given by (5) and correspond to stable solutions. In fact, the index can be interpreted as the number of eigenvalues of the eigenvalue problem
@@
The operator $\mathcal{S}$ is assumed to have eigenvalues and eigenfunctions that smoothly depend on the pseudo-arclength $\tau$. Therefore, it is not sufficient to merely determine the sign of the eigenvalue; it is also necessary to account for the additional boundary contribution. Alternatively, one can consider the (5) directly using numerical methods. Upon discretization, it leads to a Global Stiffness Matrix whose eigenvalues dictate stability - positive eigenvalues imply local minimum and thus indicate stability. The index for the case of fixed-fixed ends is shown to be equal to the number of conjugate points (for more details refer to [21]). However, we will not pursue our analysis in that direction. Instead, we track changes in the index at critical points.
@@
An eigenvalue is called a simple eigenvalue if its algebraic multiplicity is one. The meaning of (15b) becomes clearer towards the end. We analyze the behavior of the second-variational functional in the vicinity of these folds, which involves simultaneously monitoring the eigenvalues and the boundary term contribution from the corresponding eigenfunctions. We assume that changes in stability occur exclusively in the vicinity of critical points, such as folds or bifurcation points [2].
@@ 2.2 Varying Parameter at the Fixed-end
are non-zero, and (25) provides the direction in which the second variation crosses zero. The non-zero value of (26) is same as the condition (15b). The sign of $\ddot{\xi}_{0}$ indicates the direction in which the fold opens. A positive $\ddot{\xi}_{0}$ represents the fold opening towards the right, while a negative $\ddot{\xi}_{0}$ represents the fold opening toward the left. The distinguished bifurcation diagram is defined as the plot of $\left(\mathcal{L}_{\boldsymbol{\zeta}^{\prime}}\cdot\frac{\partial\boldsymbol{\zeta}_{o}}{\partial\xi_{0}}\right)\Bigg{|}_{s=0}$ vs. $\xi_{0}$ along the branches of critical points. As a simple fold is traversed, the critical eigenvalue $\mu$ and the associated eigenvector result in a sign change of $\delta^{2}J$ dictated by the signs of $\ddot{\xi}_{0}$ and $\left(\mathcal{L}_{\boldsymbol{\zeta}^{\prime}}\cdot\frac{\partial\boldsymbol{\zeta}_{o}}{\partial\xi_{0}}\right)\Bigg{|}_{s=0}$. Therefore, the index, which is the number of directions along which the second variation is negative is altered by one at each fold, and the direction of change can be found by examining the shape of the projection of the solution branch as illustrated in Figure 1(a). Indeed, the expression for the ordinate is the same as in the case of fixed-fixed ends [9], where the varying parameter appears at the fixed end $s=l$.
@@ 3.1 Setup
where $F_{x}$ and $F_{y}$ are the components of $\mathbf{F}$ along the $\mathbf{e}_{x}$ and $\mathbf{e}_{y}$ directions, respectively. In fact, the Euler-Lagrange equations given by (37) are equivalent to planar versions of the force and moment balance in elastic rods [14]
@@
In this apparatus, the parameters that describe the Dirichlet boundary conditions, such as $\theta_{o}$ and those that characterize the natural boundary conditions such as $|\mathbf{F}|$, $\Delta$, $\alpha$ and $\psi$ can be varied. We perform several numerical experiments using parameter continuation [1] and employ the proposed distinguished bifurcation diagrams to assess the stability of the equilibria.
@@ 3.2 Bifurcation Analysis
The distinguished bifurcation diagram provides information on changes in stability. So, the stability of atleast one equilibrium along a branch must be established before proceeding. For this reason, we conduct bifurcation analysis on the current apparatus and establish stability indices for some equilibria. The buckling characteristics of the elastica are well known; now, let us examine how the lever arm influences this behavior. For this analysis, we set $\alpha=0$ and $\psi=0$, and consequently we have $\lambda_{x}=0$ and $\lambda_{y}=-|\mathbf{F}|$. The classical elastica is obtained when $\Delta=0$. Before proceeding further, we non-dimensionalize (37b) by substituting $\bar{s}=\frac{s}{l}$, $\epsilon=\frac{\Delta}{l}$ and $P=\frac{\\ |\mathbf{F}|l^{2}}{K}$ that fetches
@@
The non-trivial solutions satisfying this relation that correspond to bifurcations are depicted in Figure 3. The critical buckling modes of elastica are the solutions pertaining to $\epsilon=0$ denoted as $P^{o}_{cr}=\frac{(2k-1)^{2}\pi^{2}}{4},k=1,2,\dots,\infty$. Let $P_{cr}^{+}$ represent the critical load values pertaining to positive $\epsilon$, while $P_{cr}^{-}$ represent pertaining to negative $\epsilon$. Note that for each $k$ th mode, $(k-1)^{2}\pi^{2}<P_{cr,k}^{+}<\left(\frac{2k-1}{2}\right)^{2}\pi^{2}$ and $\left(\frac{2k-1}{2}\right)^{2}\pi^{2}<P_{cr,k}^{-}<k^{2}\pi^{2}$, where $k=1,2,\dots,\infty$. It can be easily shown that these points correspond to supercritical pitchfork bifurcations, following a similar analysis in  [14, pp, 174-177]. The linearized equilibrium (42) can be rewritten as an eigenvalue problem:
@@
Typically, at a supercritical bifurcation [4], the number of negative eigenvalues increases by one with each successive bifurcation. In our setup, the index of the straight equilibria (trivial solutions) for load $P_{cr,k}^{+}<P<P_{cr,k+1}^{+}$ for positive $\epsilon$ (and $P_{cr,k}^{-}<P<P_{cr,k+1}^{-}$ for negative $\epsilon$) is $k$. For a very high $\epsilon$ value, i.e., when the $\sqrt{P}\epsilon$ line approaches the vertical line, the value of $P_{cr}^{+}$ approaches zero, suggesting that bifurcations occur at a very small load when the arm is long. On the other hand, $P_{cr}^{-}$ approaches $\pi^{2}$ for higher loads, which indicates the critical loads for the case of fixed-fixed ends. The trivial solutions are stable only for loads below $P_{cr,1}^{+}$ for positive $\epsilon$ (and below $P_{cr,1}^{-}$ for negative $\epsilon$). Consider two instances of loads which correspond to first mode and second mode, say $\pi^{2}/4$ and $9\pi^{2}/4$. These chosen straight equilibria correspond to index one and two respectively, while the negative $\epsilon$ leads to index zero and index one respectively. We use these equilibria as initial solutions for numerical continuation in the subsequent sections.

### v6

@@ 1 Introduction
The study of physical problems using the variational framework has a rich history. In this setting, the equilibria are characterized as critical points of an energy-like functional. Typically, these problems are expressed in terms of a control parameter, generating a parameter family of equilibria. Indeed, the equilibria are computed by a slow scan of a parameter from a known solution using a continuation technique [DOEDEL1991a]. However, not all equilibria are stable and physically relevant; they must correspond to the local minima of the functional. Hence, it is crucial to determine whether the critical points are local minima (or stable) and track their changes. Generally, the stability of parameter-dependent equilibria changes at singularities, such as a fold or bifurcation in the parameter space [Sattinger1972, Rabinowitz1973, iooss2012elementary]. However, it remains unclear whether a fold in the unstable equilibrium results in a stable equilibrium or a higher-order unstable equilibrium. The direction of stability transitions can be encoded through a particular projection of solutions, using distinguished bifurcation diagrams, and the stability can be assessed with a little observation and without any additional analysis. This idea of topologically exploiting the variational structure for stability examination was first employed in Astrophysics [Katz1, Katz2]. Distinguished bifurcation diagrams were later developed for finite-dimensional mechanical systems [thompson1979stability] that are subsequently extended to infinite-dimensional systems [Maddocks1987, Hoffman2005]. Since then, several versions of these diagrams have appeared in investigations of elastic problems [neukirch2002a, Heijden2003]. In most studies, the diagrams are limited to cases with fixed-fixed ends or cases where the varying parameter appears within the integrand or as a Lagrange multiplier. In this study, we generalize the distinguished bifurcation diagrams to variational problems subject to fixed-free ends, when the bifurcation parameter appears in the boundary conditions. Consequently, two relevant cases arise: when the parameter appears at the fixed end and when it appears at the free end. For this analysis, we introduce the notion of a stability index that is equal to the dimension of the subspace over which the associated second variation operator is negative. The critical points with index zero are local minima and stable. Alternatively, the energy of the system can be used to determine the direction of stability [Maddocks1995, Oreilly2011]. The lower branch of the fold corresponds to the lower-energy state and therefore, local minima.
@@
The developed framework is applied to test the stability of elastica under an end load through a lever arm. Cantilever structures with a lever arm are common in nature and engineering—for example, in plants supporting fruits or in load-bearing flexible systems. In recent years, nonlinear rod models [antman2006nonlinear] have been actively employed in the fields of biophysics [Manning_DNA], computer graphics [romero2021physical, Oreilly2012] and soft robot manipulators [Armanini2017, nayak2019]. Elastica, a simple planar version of this rod model, has equally received attention. A flexible elastic structure subjected to loads or constraints may exhibit multiple equilibria. Hence, it is crucial to determine their stability, as only stable structures are physically viable. Unstable equilibria may exist under specific constraints or boundary conditions and transitions to the available stable equilibrium under disturbances. Indeed, the presence of instability also hints at snap-back phenomena, a characteristic of multi-stable systems [Hu_2015]. Numerous studies have been conducted on the stability of elastic rods, which involve either the Jacobi test [Manning1998, Hoffman2002, Hoffman2004, Batista2015, dhanakoti2024stabilitycantileverlikestructuresapplications, Levyakov2010] or Hessian eigenvalue determination [goriely1997nonlinear, kumar2010generalized]. Alternative techniques based on bifurcation analysis are also available [Wang2021]. In this article, we assess the stability without performing any such rigorous computations. But, instead, conclusions are drawn by qualitatively examining the plot of appropriate projection of the equilibria against the parameter.
@@
Advances in material science produced highly deformable alloys and polymers, accelerating the development of soft robotics. Generally, soft robots employ very flexible structures to generate compliant mechanisms. A flexible elastic structure under load can exhibit multiple equilibria, where one equilibrium may transition to another through snap-back instability under disturbances. These bistable or multi-stable systems exhibit large rapid displacements for small stimuli when operated around the regions of instabilities. As the system surpasses the initial energy barrier, it transitions to a new state, accompanied by a release of energy. Some of nature’s quickest mechanisms seen in Mantis shrimp [patek2004deadly], Venus fly trap [skotheim2005physical], and Hummingbird beak [smith2011] can be attributed to this snap-back instability. These mechanisms are widely used in the design of high-performance soft actuators [luo2024intrinsically], soft robots [Jin2023], deployable and morphing structures [Yang2023] and haptic devices [Chen2018]. Inspired by these expanding applications, we aim to investigate elastica with a lever arm and the dependence of its stability on system parameters. These findings could offer invaluable insights for designing innovative mechanisms. Particularly, this study signifies the distinguished bifurcation diagrams for the analysis of multi-stable systems in engineering applications.
@@ 2 The Unconstrained Variational Problem
In this section, we extend the theory of distinguished bifurcation diagrams developed by [Maddocks1987] and [Hoffman2005] to variational problems subject to fixed-free ends. We consider both cases, where the bifurcation parameter appears at the fixed end and when it appears at the free end. In section 2.1, the problem and the relevant notation are introduced. The distinguished diagram at a simple fold when the parameter appears at the fixed end is developed in section 2.2 and when the parameter appears at the free end is developed in section 2.3.
@@ 2.1 Problem Formulation
while the other end $s=l$ is held free. The scalar function $\mathcal{L}$ has continuous second derivatives with respect to all its arguments and is convex in its second argument. On the other hand, the scalar function $B$ has continuous second derivatives with respect to its first argument and has continuous first derivatives with respect to the other arguments. The system can be influenced by the parameters $\xi_{0}$ and $\xi_{l}$ at the boundaries $s=0$ and $s=l$ respectively. In the current study, we vary either $\xi_{0}$ or $\xi_{l}$, but not both at a time. Maddocks [Maddocks1987] generated the bifurcation diagrams for problems when the parameter appears in the integrand and with homogeneous boundary conditions, which were later extended to the case when the parameter appears in the fixed end in a problem with non-homogeneous boundary conditions [Hoffman2005]. However, these studies are confined to cases of fixed-fixed ends. Here, we extend the analysis to the case where one of the fixed ends is set free. The first-order condition yields the critical points $\boldsymbol{\zeta}_{o}(s,\tau,\xi_{i})$, $i=0$ or $l$, of (1) as solutions to the well-known Euler-Lagrange equations
@@
The notation $\xi_{i}$ refers to either $\xi_{0}$ or $\xi_{l}$ when not explicitly specified. These critical points $\boldsymbol{\zeta}_{o}(s,\tau,\xi_{i})$ exist as a continuous curve of solutions along a branch for some pseudo-arclength parameterization $\tau$ [DOEDEL1991a] satisfying the property
@@
The notation $\dot{\boldsymbol{\zeta}}(s,\tau)$ denotes the derivative with respect to pseudo-arclength $\tau$, whereas ${\boldsymbol{\zeta}}^{\prime}(s,\tau)$ denotes the derivative with respect to independent parameter $s$. The second variation $\delta^{2}J$ evaluated at the critical points $\boldsymbol{\zeta}_{o}(s,\tau,\xi_{i})$, must satisfy the condition [gelfand1963calculus]:
@@
and $\langle\cdot,\cdot\rangle$ represents the standard inner-product in $L^{2}$- space. The coefficient $\frac{1}{2}$ has no effect on the results and is omitted in the subsequent analysis. We define an notion of index [Morse1951] which denotes the dimension of the $\mathbf{h}$ subspace over which the second variation denoted by (9) is negative. The zero index solutions satisfy the necessary condition for local minima given by (5) and correspond to stable solutions. In fact, the index can be interpreted as the number of eigenvalues of the eigenvalue problem
@@
The operator $\mathcal{S}$ is assumed to have eigenvalues and eigenfunctions that smoothly depend on the pseudo-arclength $\tau$. Therefore, it is not sufficient to merely determine the sign of the eigenvalue; it is also necessary to account for the additional boundary contribution. Alternatively, one can consider the (5) directly using numerical methods. Upon discretization, it leads to a Global Stiffness Matrix whose eigenvalues dictate stability - positive eigenvalues imply local minimum and thus indicate stability. The index for the case of fixed-fixed ends is shown to be equal to the number of conjugate points (for more details refer to [Manning1998]). However, we will not pursue our analysis in that direction. Instead, we track changes in the index at critical points.
@@
An eigenvalue is called a simple eigenvalue if its algebraic multiplicity is one. The meaning of (15b) becomes clearer towards the end. We analyze the behavior of the second-variational functional in the vicinity of these folds, which involves simultaneously monitoring the eigenvalues and the boundary term contribution from the corresponding eigenfunctions. We assume that changes in stability occur exclusively in the vicinity of critical points, such as folds or bifurcation points [Sattinger1972].
@@ 2.2 Varying Parameter at the Fixed-end
are non-zero, and (25) provides the direction in which the second variation crosses zero. The non-zero value of (26) is same as the condition (15b). The sign of $\ddot{\xi}_{0}$ indicates the direction in which the fold opens. A positive $\ddot{\xi}_{0}$ represents the fold opening towards the right, while a negative $\ddot{\xi}_{0}$ represents the fold opening toward the left. The distinguished bifurcation diagram is defined as the plot of $\left(\mathcal{L}_{\boldsymbol{\zeta}^{\prime}}\cdot\frac{\partial\boldsymbol{\zeta}_{o}}{\partial\xi_{0}}\right)\Bigg|_{s=0}$ vs. $\xi_{0}$ along the branches of critical points. As a simple fold is traversed, the critical eigenvalue $\mu$ and the associated eigenvector result in a sign change of $\delta^{2}J$ dictated by the signs of $\ddot{\xi}_{0}$ and $\left(\mathcal{L}_{\boldsymbol{\zeta}^{\prime}}\cdot\frac{\partial\boldsymbol{\zeta}_{o}}{\partial\xi_{0}}\right)\Bigg|_{s=0}$. Therefore, the index, which is the number of directions along which the second variation is negative is altered by one at each fold, and the direction of change can be found by examining the shape of the projection of the solution branch as illustrated in Figure 1(a). Indeed, the expression for the ordinate is the same as in the case of fixed-fixed ends [Hoffman2005], where the varying parameter appears at the fixed end $s=l$.
@@ 3.1 Setup
where $F_{x}$ and $F_{y}$ are the components of $\mathbf{F}$ along the $\mathbf{e}_{x}$ and $\mathbf{e}_{y}$ directions, respectively. In fact, the Euler-Lagrange equations given by (37) are equivalent to planar versions of the force and moment balance in elastic rods [antman2006nonlinear]
@@
In this apparatus, the parameters that describe the Dirichlet boundary conditions, such as $\theta_{o}$ and those that characterize the natural boundary conditions such as $|\mathbf{F}|$, $\Delta$, $\alpha$ and $\psi$ can be varied. We perform several numerical experiments using parameter continuation [DOEDEL1991a] and employ the proposed distinguished bifurcation diagrams to assess the stability of the equilibria.
@@ 3.2 Bifurcation Analysis
The distinguished bifurcation diagram provides information on changes in stability. So, the stability of atleast one equilibrium along a branch must be established before proceeding. For this reason, we conduct bifurcation analysis on the current apparatus and establish stability indices for some equilibria. The buckling characteristics of the elastica are well known; now, let us examine how the lever arm influences this behavior. For this analysis, we set $\alpha=0$ and $\psi=0$, and consequently we have $n_{x}=0$ and $n_{y}=-|\mathbf{F}|$. The classical elastica is obtained when $\Delta=0$. Before proceeding further, we non-dimensionalize (37b) by substituting $\bar{s}=\frac{s}{l}$, $\epsilon=\frac{\Delta}{l}$ and $P=\frac{\\ |\mathbf{F}|l^{2}}{K}$ that fetches
@@
The non-trivial solutions satisfying this relation that correspond to bifurcations are depicted in Figure 3. The critical buckling modes of elastica are the solutions pertaining to $\epsilon=0$ denoted as $P^{o}_{cr}=\frac{(2k-1)^{2}\pi^{2}}{4},k=1,2,\dots,\infty$. Let $P_{cr}^{+}$ represent the critical load values pertaining to positive $\epsilon$, while $P_{cr}^{-}$ represent pertaining to negative $\epsilon$. Note that for each $k$ th mode, $(k-1)^{2}\pi^{2}<P_{cr,k}^{+}<\left(\frac{2k-1}{2}\right)^{2}\pi^{2}$ and $\left(\frac{2k-1}{2}\right)^{2}\pi^{2}<P_{cr,k}^{-}<k^{2}\pi^{2}$, where $k=1,2,\dots,\infty$. It can be easily shown that these points correspond to supercritical pitchfork bifurcations, following a similar analysis in  [antman2006nonlinear, pp, 174-177]. The linearized equilibrium (42) can be rewritten as an eigenvalue problem:
@@
Typically, at a supercritical bifurcation [iooss2012elementary], the number of negative eigenvalues increases by one with each successive bifurcation. In our setup, the index of the straight equilibria (trivial solutions) for load $P_{cr,k}^{+}<P<P_{cr,k+1}^{+}$ for positive $\epsilon$ (and $P_{cr,k}^{-}<P<P_{cr,k+1}^{-}$ for negative $\epsilon$) is $k$. For a very high $\epsilon$ value, i.e., when the $\sqrt{P}\epsilon$ line approaches the vertical line, the value of $P_{cr}^{+}$ approaches zero, suggesting that bifurcations occur at a very small load when the arm is long. On the other hand, $P_{cr}^{-}$ approaches $\pi^{2}$ for higher loads, which indicates the critical loads for the case of fixed-fixed ends. The trivial solutions are stable only for loads below $P_{cr,1}^{+}$ for positive $\epsilon$ (and below $P_{cr,1}^{-}$ for negative $\epsilon$). Consider two instances of loads which correspond to first mode and second mode, say $\pi^{2}/4$ and $9\pi^{2}/4$. These chosen straight equilibria correspond to index one and two respectively, while the negative $\epsilon$ leads to index zero and index one respectively. We use these equilibria as initial solutions for numerical continuation in the subsequent sections.